tan[ 1/2 sin^-1 (2a/1+a^2) + 1/2 cos^-1 (1-a^2/1+a^2) ]

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

tan[ 1/2 sin^-1 (2a/1+a^2) + 1/2 cos^-1 (1-a^2/1+a^2) ]=tan[tan-1a +tan-1a]=tan(2tan-1a)=2a/(1-a2)

Godfrey Classic Prince
12 years ago

Dear Nelwin Jacob,

tan[ 1/2 sin^-1 (2a/1+a^2) + 1/2 cos^-1 (1-a^2/1+a^2) ] = tan[tan-1a +tan-1a]=tan(2tan-1a)=2a/(1-a2)

 

Hope this helped you immensely...

All the Very Best & Good Luck to you ...

Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-Madras

Please approve my answer if you liked it by clicking on "Yes" given below...!!Smile